An electron is accelerated through 2 400 V from rest and then enters a uniform 1.70-T magnetic field. What are (a) the maximum and (b) the minimum values of the magnetic force this particle can experience?


1
Expert's answer
2021-02-08T18:42:03-0500

a)



"qU=0.5mv^2\\v=\sqrt{\frac{2qU}{m}}\\F=qvB
\\F=qB\sqrt{\frac{2qU}{m}}\\F=(1.6\cdot10^{-19})(1.7)\sqrt{\frac{2(2400)(1.6\cdot10^{-19})}{(9.1\cdot10^{-31})}}\\=7.9\cdot10^{-12}\ N"

b) If the magnetic field is parallel to the motion of the electron:



"F=0\ N"